Please help me solve this equation: Three times a first number decreased by a second number is 1. The first number increased by twice the second number is 12. Find the numbers

I tried: Let x equal 1st number
Let y equal 2nd number
3x-y=1
3x+2y=12

Answer by checkley77(12844)

3x-y=1 multiply by 2 & add.
x+2y=12
6x-2y=2
------------------
7x=14
x=14/7
x=2 ans.
2+2y=12
2y=12-2
2y=10
y=10/2
y=5 ans.
Proof:
3*2-5=1
6-5=1
1=1